OpenStudy (anonymous):

got statistic test tonight, help ! a sample space is composed of 3 outcomes , A,B and C .Outcome B is twice as probable as A and C is twice as probable as B. Find the probabilities of the events of A,B and C

OpenStudy (anonymous):

remember in probabilities, the sum should always be 1

OpenStudy (anonymous):

so you have A + B +C = 1 now just express A and C in terms of B then solve

OpenStudy (anonymous):

but im confuse.. what does it mean by Outcome B is twice as probable as A and C is twice as probable as B.???

OpenStudy (caozeyuan):

P(B)=2*P(A)=1/2 * P(C) @Ika_Kim

OpenStudy (caozeyuan):

so 1/2*P(B)+P(B)+2*P(B)=1, now you can find your P(B) and hence the others

OpenStudy (caozeyuan):

does my response worth a medal?

OpenStudy (anonymous):

urm...i dont understand...sorry....

OpenStudy (caozeyuan):

As a said, B is twice as possible as A literally translates to P(B)=2*P(A), do you understand this?

OpenStudy (caozeyuan):

and based on the idea above, we also have P(C) =2*P(B),

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

then ???

OpenStudy (caozeyuan):

You know that P(A)+P(B)+P(C)=1, do you?

OpenStudy (anonymous):

yes i do.. so... we convert them to the same term ?

OpenStudy (caozeyuan):

so you can treat P(A),P(B),P(C) as three variables and the three equation appeared before are the ones to solve, and I assume you can solve a system of linear equations with three variables

OpenStudy (caozeyuan):

thx @xlegendx , you just boosted my score by 1

OpenStudy (anonymous):

owhh ok. i got it !!!! thanks guys !!!!!!!! :D